OpenStudy (anonymous):

Not sure how to do this one...it says:Evaluate the limit lim x approaches -6 (1/6+1/x)/ 6+x

OpenStudy (anonymous):

jus substitute -6 into the function ,,,lol

OpenStudy (kirbykirby):

Try putting the numerator terms on the same denominator. You should get \(\dfrac{1}{6}+\dfrac{1}{x}=\dfrac{x+6}{6x}\) . Now notice how the denominator can be cancelled out now ;)